Solve the quadratic equation below by factoring and applying the zero product principle. (3x+4)² - 25=0
Answer by ankor@dixie-net.com(22740)

(3x+4)² - 25=0
FOIL (3x+4)(3x+4)
9x^2 + 12x + 12x + 16 - 25 = 0
9x^2 + 24x - 9 = 0
factor out 3
3(3x^2 + 8x - 3) = 0
divide both sides by 3
3x^2 + 8x - 3 = 0
Factor again
(3x - 1)(x + 3) = 0
two solutions
3x = 1
x = 1/3
and
x = -3